OpenStudy (melody323):

Which of the points on the number line shows the best approximation for the square root of 15 ? A B C D

OpenStudy (mathstudent55):

What is \(\sqrt {16} \) ?

OpenStudy (melody323):

4

OpenStudy (melody323):

@mathstudent55

Vocaloid (vocaloid):

good, that means the square root of 15 must be a little bit less than 4

Vocaloid (vocaloid):

which point matches this description?

OpenStudy (mathstudent55):

Good. What is \(\sqrt 9\) ?

OpenStudy (melody323):

its 3.87

OpenStudy (melody323):

for the square rrot oof 15

OpenStudy (melody323):

root of

Vocaloid (vocaloid):

yes,so which point on the graph matches?

OpenStudy (melody323):

the square root of 9 is 3

OpenStudy (mathstudent55):

Since \(\sqrt {16} = 4\), and \(\sqrt {9} = 3\), you expect the \(\sqrt {15}\) to be between 3 and 4. Also, since 15 is much closer to 16 than to 9, you expect \(\sqrt{15} \) to be much closer to 4 than to 3. Which of the numbers shown on the number line is between 3 and 4, but much closer to 4 than to 3?

OpenStudy (melody323):

A

OpenStudy (mathstudent55):

Correct. That is how you figure out the answer without having to calculate \(\sqrt {15}\) on a calculator.

OpenStudy (melody323):

ok thanks.

OpenStudy (mathstudent55):

You're welcome.